A flask contains 5 moles of an ideal gas at 25C and 2 atm pressure. The temperature of the gas is raised to 50C. What fraction of the gas must be removed to keep the pressure in the flask constant.
Answer #1

Initial Initial molee of ideal gas, na = 5 moles temperature, To 25°C = 25+273 = 298K Final temperature, T2 = 50°c = 50+273 =